the sum of two numbers is eighteen. the total of three times the smaller and twice the larger is forty-four. find the two numbers
Answer by checkley75(3666)

X+Y=18, X=18-Y
3X+2Y=44 NOW SUBSTITUTE (18-Y) FOR X
3(18-Y)+2Y=44
54X-3Y+2Y=44
-Y=44-54
-Y=-10
Y=10 ANSWER
X+10=18
X=18-10
X=8 ANSWER.
PROOF
3*8+2*10=44
24+20=44
44=44